OpenStudy (anonymous):

Allison has a poster that is 15 in by 18 in. What will the dimensions of the poster be if she scales it down by a factor of one-third ? 15 in by 18 in 45 in by 54 in 2.5 in by 2 in 5 in by 6 in

OpenStudy (mokeira):

\[15\times \frac{ 1 }{3 }=5\] \[18 \times \frac{ 1 }{3 }=6\] so 5in by 6in